# Impact of in-situ nuclear networks and atomic opacities on neutron star merger ejecta dynamics, nucleosynthesis, and kilonovae

### v1

@@ Abstract
Methods. Ejecta profiles from long-term numerical-relativity simulations are evolved to $\sim 30$ days using a 2D ray-by-ray approach. We compare simplified heating-rate and thermalization prescriptions with in-situ nuclear reaction network (NN) calculations that track nuclear energy deposition and include a composition-dependent thermalization scheme. We also contrast various gray opacity models with a frequency-dependent treatment based on atomic calculations.
@@
Conclusions. Coupling hydrodynamics to an in-situ NN is crucial for reliable nucleosynthesis and kilonova predictions. Resolving the first several hundreds milliseconds of the hydrodynamics is essential for robust nucleosynthesis calculations. Composition-dependent thermalization and frequency-dependent, atomic-physics-based opacities are needed to accurately capture the temperature evolution of the ejecta and the brightness and color evolution of the kilonova. Calibrated analytic nuclear-power fits with simplified thermalization and opacity prescriptions can still reproduce the density and temperature evolution of the ejecta.
@@ 1 Introduction
Binary neutron star mergers (BNSMs) are one of the main sources for the production of heavy elements via the rapid neutron capture ($r$-process) nucleosynthesis (Eichler et al., 1989; Li & Paczynski, 1998; Metzger et al., 2010; Cowan et al., 2021; Perego et al., 2021). The first multimessenger observation of a BSNM was triggered by the gravitational wave (GW) signal GW170817, and associated with the short gamma-ray burst (GRB) GRB 170817A, neutrino emissions and the kilonova AT2017gfo, i.e. the electromagnetic (EM) counterpart powered by the decay of the freshly synthesized, unstable nuclei (Abbott et al., 2017b, a, c; Coulter et al., 2017; Drout et al., 2017; Savchenko et al., 2017). Kilonova candidates have also been identified in association with other short GRBs (Berger et al., 2013; Gompertz et al., 2018; Rossi et al., 2020) and hybrid or long GRBs (Rastinejad et al., 2022; Troja et al., 2022).
@@
The properties of the matter ejected from a BNSM (mass, temperature, velocity, electron fraction) are determined by the merger and post-merger dynamics, neutrino interactions, and the nuclear EOS (Rosswog et al., 2014; Radice et al., 2018; Perego et al., 2019; Foucart et al., 2020; Bernuzzi, 2020; Fujibayashi et al., 2020; Just et al., 2022; Kiuchi et al., 2023; Lund et al., 2024; Cheong et al., 2025). The nuclear reactions burning in the expanding material depend on these initial conditions and on the thermodynamic evolution of the outflow. The latter is dominated by its hydrodynamic expansion and affected by the released nuclear energy and radiative transport effects (Korobkin et al., 2012; Lippuner & Roberts, 2015; Cowan et al., 2021; Fryer et al., 2024; Magistrelli et al., 2024). The final EM (optical, UV, near-infrared and $\gamma$) emission is dictated by the thermal energy of the ejecta, the nuclear power released in the reacting material, the way the products of the nuclear reactions thermalize with the surrounding matter, the ejecta morphology, and the atomic opacities in the outflow (Barnes et al., 2016; Kasen et al., 2017; Wollaeger et al., 2018; Kasen & Barnes, 2019; Hotokezaka & Nakar, 2019; Tanaka et al., 2020; Korobkin et al., 2019; Fontes et al., 2020; Korobkin et al., 2021; Gillanders et al., 2022; Fontes et al., 2023; Collins et al., 2023; Shingles et al., 2023; Pognan et al., 2024; Sneppen et al., 2024). To understand the nucleosynthesis in BNSM ejecta, how they can participate in the galactic enrichment, and the associated EM counterparts, it is thus necessary to model all of these complex aspects of the merger and post-merger dynamics.
@@
General-relativistic (magneto)-hydrodynamics (GRMHD) simulations implementing candidate nuclear EOSs can follow a binary system from the inspiral phase until hundreds of milliseconds after merger, collecting information on the thermodynamic properties of the outflows (e.g. Radice et al., 2016; Bauswein et al., 2020; Nedora et al., 2021; Combi & Siegel, 2023; Kiuchi et al., 2024; Fields et al., 2025). New moment-based neutrino-transport schemes and Monte Carlo methods have improved our understanding of the role of neutrinos in driving mass ejection (in particular for long-lived remnants) and setting the electron fraction, and thus initial composition, of these outflows (Foucart et al., 2016; Radice et al., 2022; Zappa et al., 2023; Kawaguchi et al., 2025). Usually, ejecta properties extracted from numerical relativity (NR) simulations are transferred to other codes for their long-term ($\sim$ days) evolution with the use of tracers or some suitable mapping (Korobkin et al., 2012; Martin et al., 2015; Barnes et al., 2016; Radice et al., 2018; Perego et al., 2022; Curtis et al., 2024; Groenewegen et al., 2025).
@@
As the material expands and cools, the temperature eventually drops below the nuclear statical equilibrium (NSE) threshold. Beyond this point, an additional source of uncertainty in modeling $r$-process nucleosynthesis arises from the poorly constrained nuclear properties of isotopes far from stability. In particular, simulating ejecta form BNSM requires knowledge of the properties of nuclei near the neutron drip line. Despite significant theoretical and experimental effort, key quantities such as nuclear masses, $\beta$-decay rates, neutron-capture cross sections, and fission yields remain highly uncertain, yet strongly impact nucleosynthesis pathways and influence EM observables (Mumpower et al., 2016; Barnes et al., 2021; Zhu et al., 2022; Martinez-Pinedo & Langanke, 2023; Mumpower et al., 2024; Martinet & Goriely, 2025).
@@
Most of the current ejecta models treat nuclear physics and radiation-hydrodynamic separately. For example, Rosswog et al. (2014); Kawaguchi et al. (2021); Wu et al. (2022); Kawaguchi et al. (2024) input pre-computed, analytical nuclear powers fits in the energy equation. In Korobkin et al. (2012); Radice et al. (2018); Perego et al. (2022); Gillanders et al. (2022); Just et al. (2022); Curtis et al. (2024), the nucleosynthesis is calculated via a post-processing method by prescribing a simplified homologous expansion and feeding it to a nuclear reaction network (NN). Similar simplifications have been used to perform full Monte Carlo radiative-transport simulations to calculate kilonova light curves and spectra (Collins et al., 2023; Shingles et al., 2023; Pognan et al., 2024). These methods neglect the complex interplay between fluid dynamics, nucleosynthesis, and radiative transport, and fail to capture the nonlinear feedback of radioactive heating and composition-dependent radiation-matter interactions on the outflow dynamics. To address this limitation and study the interplay between ejecta dynamics and nucleosynthesis, Magistrelli et al. (2024) performed the first analysis of BNSM ejecta evolution that couples ray-by-ray radiation-hydrodynamic expansion with an in-situ NN. This more self-consistent modeling revealed qualitative differences in the predicted kilonova emission and nucleosynthesis. The same framework was applied in Bernuzzi et al. (2025) and Jacobi et al. (2025) to investigate outflows from asymmetric BNSMs.
@@
In this work, we assess the impact of in-situ nuclear burning models and atomic opacities in radiation-hydrodynamics BNSM ejecta simulations. In order to explore efficiently different models, we focus on a 2D ray-by-ray setup where radiation-hydrodynamics is treated in the diffusion limit. In Sec. 2, we describe our frequency-dependent radiation-hydrodynamics setup and its coupling with the in-situ NN, which updates the kNECnn code (Magistrelli et al., 2024, see also Morozova et al. (2015); Wu et al. (2022)). The code implements simple, commonly used analytical prescriptions for the nuclear burning and opacities as well as different models for composition-informed thermalization and opacities enabled by the NN coupling. We also discuss the extraction of initial ejecta profiles from NR simulations and describe how we incorporate the effects of a jet depositing energy into the polar regions of the ejecta. Further details on the frequency-dependent radiative transport equations and on the initialization of the NN can be found in Appendices A and B. In Sec. 3, we analyze the impact of each physical improvement introduced here relative to previous works. We examine how the inclusion of an in-situ NN alters nucleosynthesis and light-curve predictions, and we asses the improvements enabled by the resulting composition-dependent thermalization and opacity models. We also study the impact of the additional energy deposited by a polar jet. We summarize our findings in Sec. 4.
@@ 2.1 Ray-by-ray Radiation-Hydrodynamics
The ejecta are evolved by solving the system of Lagrangian radiation-hydrodynamic equations presented in SNEC (Morozova et al., 2015). Non-spherical effects in the dynamics are approximately incorporated under the axisymmetry and ray-by-ray assumptions. The polar angle dependency of the ejecta properties is taken into account by mapping each angular section into an effective 1D, spherically symmetric problem. The total mass of each section is scaled by the factor $\lambda_{\theta}=4\pi/\Delta\Omega$, where $\Delta\Omega$ is the solid angle subtended by the angular section. The sections are evolved independently, and the hydrodynamic variables and nucleosynthesis results are then recombined via mass weighted averages. The mapping procedure assures that the intensive quantities are always preserved (Magistrelli et al., 2024). Note that non-radial flows of matter and radiation are neglected, as well as higher dimensional fluid instabilities (e.g. Kelvin-Helmholtz), convection and the possibility of shells surpassing each other. These assumptions could in principle lead to overestimating the pressure work exchanged between the radial shells and introduce artificial shocks in the system.
@@
where the radiative flux for each shell is calculated under the assumption of local thermodynamical equilibrium (LTE) as
@@
proportional to the one defined in Levermore & Pomraning (1981) (see Appendix A for further details).
@@ 2.2 Nuclear Network Coupling
Within every hydrodynamic time-step $\Delta t$, each fluid element independently evolves the associated NN, which automatically defines its own sub-steps based on the nuclear timescales. Let’s call $t_{0}$ the initial time of the hydrodynamical time-step. During the NN evolution, the density is prescribed by a log-interpolation between $\rho(t_{0})$ and $\rho(t_{0}+\Delta t)$. The latter is calculated from $\rho(t_{0})$ and the velocity of the shell boundaries at $t=t_{0}$, updated via the momentum transport equation (for which no nuclear input is required). During the timestep, we assume a constant $T=T(t_{0})$ for the NN calculations (no self-heating). The temperature will then be updated by the second part of the hydrodynamic step. We collect the energy released by nuclear reactions during each sub-step, and define the (non-thermalized) nuclear power $\dot{q}_{\text{nucl}}(t_{0})$ as the total energy produced, divided by $\Delta t$. This quantity will enter the energy equation (1) after being thermalized into $\dot{\epsilon}_{\text{nucl}}(t_{0})$ as described in Sec. 2.4. At the end of the NN calculations, we use the updated isotopic composition to get the mean molecular weight, which enters the EOS, and the opacity profile, needed to calculate the luminosity term in Eq. (1) and the observed fluxes. In other words, we use the abundances at $t=t_{0}+\Delta t$ as a representation of the composition during the whole hydrodynamic step. This procedure produces all the information needed to evolve Eq. (1), and thus to update temperature and velocity of the outflow.
@@
For matter that experienced hot NSE during the NR simulation, but underwent NSE freeze-out before reaching the extraction radius, the Boltzmann term in Eq. (8) tends to overestimate the abundances of the iron-group elements. Moreover, the impact of nuclear reactions happening immediately after NSE drop-out is neglected. To correct these effects, we implement a backtracking-NSE initialization similar to the one described in Radice et al. (2016); Perego et al. (2022). Essentially, if $T_{0}$, $\rho_{0}$, $Y_{e,0}$, $s_{0}$ are the properties at the extraction radius, we assume that the fluid element had expanded adiabatically from its NSE configuration at some $T_{\text{NSE}}>T_{0}$, $\rho_{\text{NSE}}$, $Y_{e,\text{NSE}}$, $s_{\text{NSE}}$. On the small timescales ($\lesssim 10$ ms) needed for the ejected matter to get to the extraction sphere, we expect the electron fraction and the entropy not to change significantly, so $Y_{e,\text{NSE}}\simeq Y_{e,0}$, $s_{\text{NSE}}\simeq s_{0}$. We check a posteriori that the relative changes in the entropy and electron fraction are of the order of 5% or less during the pre-evolution phase in most part of the ejecta (see Fig. 19). We finally estimate the density $\rho_{\text{NSE}}$ of the fluid element at NSE freeze-out by using the EOS employed in the NR simulation and starting from $T_{\text{NSE}}$, $Y_{e,\text{NSE}}=Y_{e,0}$, $s_{\text{NSE}}=s_{0}$. We calculate the correspondent NSE composition via Eq. (8), and then mimic the abundance evolution up to the extraction radius assuming homologous expansion. In particular, we parametrize the density evolution as in Eq. (1) of Lippuner & Roberts (2015), and estimate the expansion timescale $\tau_{\text{ex}}$ as in Radice et al. (2016) [^1] The prescription of Lippuner & Roberts (2015) is meant to approximate the physical evolution of a fluid element starting at early times after the ejection. The assumption of homologous expansion, not realistic at early times, is needed to get an estimate of the expansion timescale of the ejecta. Since variations of a factor of a few in $\tau_{\text{ex}}$ do not significantly affect the evolution of the NN (Lippuner & Roberts, 2015; Perego et al., 2021), our assumption gives a reliable approximation for the composition at the extraction radius.. To ensure that the evolved thermodynamic trajectory matches the physical conditions of the fluid element at the extraction radius, we also use an analytical prescription for the temperature history inspired by the adiabatic expansion law for an ideal gas, $T(\rho)=T_{\text{NSE}}\,(\rho(t)/\rho_{\text{NSE}})^{(\gamma-1)}$. The effective adiabatic factor $\gamma$, specific for each fluid element, is determined by $\gamma=1+\ln(\rho_{0}/\rho_{\text{NSE}})/\ln(T_{0}/T_{\text{NSE}})$, which ensures $T(\rho_{0})=T_{0}$. The composition at the end of the backtracking is used together with the originally extracted $s$ to define the initial condition of the fluid element.
@@
Low-entropy tidal ejecta (more efficiently produced for strongly asymmetric binaries) are expelled before the merger can raise the temperature, and are therefore expected to retain a composition close to the one set by the cold-NSE condition, for which $T\lesssim 1$ MeV. In principle, this would suggest using the cold-NSE method. However, in an NR simulation, numerical artifacts (e.g. shock heating propagating inwards from the surface of the compact objects) can boost the initial temperature of the neutron stars up to few MeV. In all the NR simulations analyzed in this paper, the neutron stars reach this temperature in $\sim 1$ ms during the inspiral phase. The spurious high temperatures introduce small artifacts in the tracers calculations and in the initialized NSE composition, regardless of the method used. In Appendix B, we investigate the systematic uncertainties propagating from the different initialization methods to the nucleosynthesis and kilonova results.

### v2

# Impact of in situ nuclear networks and atomic opacities on neutron star merger ejecta dynamics, nucleosynthesis, and kilonovae

@@ Abstract
Methods. Ejecta profiles from long-term numerical-relativity simulations of asymmetric neutron star binaries with a massive neutron star remnant are evolved to $\sim 30$ days using a 2D ray-by-ray approach. We compare simplified heating-rate and thermalization prescriptions with in-situ Nuclear reaction Network (NN) calculations that track nuclear energy deposition and include a composition-dependent thermalization scheme. We also contrast various gray opacity models with a frequency-dependent treatment based on atomic calculations.
@@
Conclusions. Coupling hydrodynamics to an in-situ NN is crucial for reliable nucleosynthesis and kilonova predictions. Resolving the first several hundred milliseconds of the hydrodynamics is essential for robust nucleosynthesis calculations. Composition-dependent thermalization and frequency-dependent, atomic-physics-based opacities are needed to accurately capture the temperature evolution of the ejecta and the brightness and color evolution of the kilonova. Calibrated analytic nuclear-power fits with simplified thermalization and opacity prescriptions can still reproduce the density and temperature evolution of the ejecta.
@@ 1 Introduction
Binary neutron star mergers (BNSMs) are one of the main sources for the production of heavy elements via the rapid neutron capture ($r$-process) nucleosynthesis (Eichler et al., 1989; Li & Paczynski, 1998; Metzger et al., 2010; Cowan et al., 2021; Perego et al., 2021). The first multi-messenger observation of a BSNM was triggered by the gravitational wave (GW) signal GW170817, and associated with the short gamma-ray burst (GRB) GRB 170817A and the kilonova AT2017gfo, i.e. the electromagnetic (EM) counterpart powered by the decay of the freshly synthesized, unstable nuclei (Abbott et al., 2017b, a, c; Coulter et al., 2017; Drout et al., 2017; Savchenko et al., 2017). Kilonova candidates have also been identified in association with other short GRBs (Berger et al., 2013; Gompertz et al., 2018; Rossi et al., 2020) and hybrid or long GRBs (Rastinejad et al., 2022; Troja et al., 2022).
@@
The properties of the matter ejected from a BNSM (mass, temperature, velocity, and electron fraction) are determined by the merger and post-merger dynamics, neutrino interactions, and the nuclear equation of state (EOS) (Rosswog et al., 2014; Radice et al., 2018; Perego et al., 2019; Foucart et al., 2020; Bernuzzi, 2020; Fujibayashi et al., 2020; Just et al., 2022; Kiuchi et al., 2023; Lund et al., 2024; Cheong et al., 2025). The nuclear reactions burning in the expanding material depend on these initial conditions and on the thermodynamic evolution of the outflow. The latter is dominated by its hydrodynamic expansion and affected by the released nuclear energy and radiative transport effects (Korobkin et al., 2012; Lippuner & Roberts, 2015; Cowan et al., 2021; Fryer et al., 2024; Magistrelli et al., 2024). The final EM (optical, UV, near-infrared, and $\gamma$) emission is dictated by the thermal energy of the ejecta, the nuclear power released in the reacting material, the way the products of the nuclear reactions thermalize with the surrounding matter, the ejecta morphology, and the atomic opacities in the outflow (Barnes et al., 2016; Kasen et al., 2017; Wollaeger et al., 2018; Kasen & Barnes, 2019; Hotokezaka & Nakar, 2019; Tanaka et al., 2020; Korobkin et al., 2019; Fontes et al., 2020; Korobkin et al., 2021; Gillanders et al., 2022; Fontes et al., 2023; Collins et al., 2023; Shingles et al., 2023; Pognan et al., 2024; Sneppen et al., 2024). To understand the nucleosynthesis in BNSM ejecta, how they can participate in the galactic enrichment, and the associated EM counterparts, it is thus necessary to model all of these complex aspects of the merger and post-merger dynamics.
@@
General-relativistic (magneto)-hydrodynamics (GRMHD) simulations implementing candidate nuclear EOSs can follow a binary system from the inspiral phase until hundreds of milliseconds after merger, collecting information on the thermodynamic properties of the outflows (e.g. Radice et al., 2016; Bauswein et al., 2020; Nedora et al., 2021; Combi & Siegel, 2023; Kiuchi et al., 2024; Fields et al., 2025). New moment-based neutrino-transport schemes and Monte Carlo methods have improved our understanding of the role of neutrinos in driving mass ejection (in particular for long-lived remnants) and setting the electron fraction, and thus initial composition, of these outflows (Foucart et al., 2016; Radice et al., 2022; Zappa et al., 2023; Kawaguchi et al., 2025). Usually, ejecta properties extracted from numerical relativity (NR) simulations are transferred to other codes for their long-term (approximately days) evolution with the use of tracers or some suitable mapping (Korobkin et al., 2012; Martin et al., 2015; Barnes et al., 2016; Radice et al., 2018; Perego et al., 2022; Curtis et al., 2024; Groenewegen et al., 2025).
@@
As the material expands and cools, the temperature eventually drops below the nuclear statical equilibrium (NSE) threshold. Beyond this point, an additional source of uncertainty in modeling $r$-process nucleosynthesis arises from the poorly constrained nuclear properties of isotopes far from stability. In particular, simulating ejecta from BNSMs requires knowledge of the properties of nuclei near the neutron drip line. Despite significant theoretical and experimental effort, key quantities such as nuclear masses, $\beta$-decay rates, neutron-capture cross sections, and fission yields remain highly uncertain, yet they strongly impact nucleosynthesis pathways and influence EM observables (Mumpower et al., 2016; Barnes et al., 2021; Zhu et al., 2022; Martinez-Pinedo & Langanke, 2023; Mumpower et al., 2024; Martinet & Goriely, 2025).
@@
Most of the current ejecta models treat nuclear physics and radiation-hydrodynamic separately. For example, Rosswog et al. (2014); Kawaguchi et al. (2021); Wu et al. (2022); Kawaguchi et al. (2024) input precomputed, analytical nuclear powers fits in the energy equation. In Korobkin et al. (2012); Radice et al. (2018); Perego et al. (2022); Gillanders et al. (2022); Just et al. (2022); Curtis et al. (2024), the nucleosynthesis is calculated via a post-processing method by prescribing a simplified homologous expansion and feeding it to a Nuclear reaction Network (NN). Similar simplifications have been used to perform full Monte Carlo radiative-transport simulations to calculate kilonova light curves and spectra (Collins et al., 2023; Shingles et al., 2023; Pognan et al., 2024). These methods neglect the complex interplay between fluid dynamics, nucleosynthesis, and radiative transport, and they fail to capture the nonlinear feedback of radioactive heating and composition-dependent radiation-matter interactions on the outflow dynamics. To address this limitation and study the interplay between ejecta dynamics and nucleosynthesis, Magistrelli et al. (2024) performed the first analysis of BNSM ejecta evolution that couples ray-by-ray radiation-hydrodynamic expansion with an in-situ NN. This more self-consistent modeling revealed qualitative differences in the predicted kilonova emission and nucleosynthesis. The same framework was applied in Bernuzzi et al. (2025) and Jacobi et al. (2026) to investigate outflows from asymmetric BNSMs.
@@
In this work, we assess the impact of in-situ nuclear-burning models and atomic opacities in radiation-hydrodynamics BNSM ejecta simulations. In order to explore efficiently different models, we focus on a 2D ray-by-ray setup where radiation hydrodynamics is treated in the diffusion limit. In Sec. 2, we describe our frequency-dependent radiation-hydrodynamics setup and its coupling with the in-situ NN, which updates the kNECnn code (Magistrelli et al., 2024, see also Morozova et al. (2015); Wu et al. (2022)). The code implements simple, commonly used analytical prescriptions for the nuclear burning and opacities as well as different models for composition-informed thermalization and opacities enabled by the NN coupling. We also discuss the extraction of initial ejecta profiles from NR simulations and describe how we incorporate the effects of a jet depositing energy into the polar regions of the ejecta. Further details on the frequency-dependent radiative transport equations and on the initialization of the NN can be found in Appendices A and B. In Sec. 3, we analyze the impact of each physical improvement introduced here relative to previous works. We examine how the inclusion of an in-situ NN alters nucleosynthesis and light-curve predictions, and we asses the improvements enabled by the resulting composition-dependent thermalization and opacity models. We also study the impact of the additional energy deposited by a polar jet. We summarize our findings in Sec. 4.
@@ 2.1 Ray-by-ray radiation-hydrodynamics
The ejecta are evolved by solving the system of Lagrangian radiation-hydrodynamic equations presented in SNEC (Morozova et al., 2015). Nonspherical effects in the dynamics are approximately incorporated under the axisymmetry and ray-by-ray assumptions. The polar angle dependency of the ejecta properties is taken into account by mapping each angular section into an effective 1D, spherically symmetric problem. The total mass of each section is scaled by the factor $\lambda_{\theta}=4\pi/\Delta\Omega$, where $\Delta\Omega$ is the solid angle subtended by the angular section. The sections are evolved independently, and the hydrodynamic variables and nucleosynthesis results are then recombined via mass weighted averages. The mapping procedure assures that the intensive quantities are always preserved (Magistrelli et al., 2024). Note that non-radial flows of matter and radiation are neglected, as well as higher dimensional fluid instabilities (e.g. Kelvin-Helmholtz), convection and the possibility of shells surpassing each other. These assumptions could in principle lead to overestimating the pressure work exchanged between the radial shells and introduce artificial shocks in the system.
@@
where the radiative flux for each shell is calculated under the assumption of local thermodynamical equilibrium (LTE) and in the single-temperature diffusion approximation as
@@
which is proportional to the one defined in Levermore & Pomraning (1981) (see Appendix A for further details).
@@ 2.2 Nuclear network coupling
Within every hydrodynamic time-step $\Delta t$, each fluid element independently evolves the associated NN, which automatically defines its own sub-steps based on the nuclear timescales. We call $t_{0}$ the initial time of the hydrodynamical time-step. During the NN evolution, the density is prescribed by a log-interpolation between $\rho(t_{0})$ and $\rho(t_{0}+\Delta t)$. The latter is calculated from $\rho(t_{0})$ and the velocity of the shell boundaries updated via the momentum transport equation (for which no nuclear input is required). During the timestep, we assume a constant $T=T(t_{0})$ for the NN calculations (no self-heating). The temperature will then be updated by the second part of the hydrodynamic step. We collect the energy released by nuclear reactions during each sub-step, and define the (non-thermalized) nuclear power $\dot{q}_{\text{nucl}}(t_{0})$ as the total energy produced, divided by $\Delta t$. This quantity will enter the energy equation (1) after being thermalized into $\dot{\epsilon}_{\text{nucl}}(t_{0})$ as described in Sec. 2.4. At the end of the NN calculations, we use the updated isotopic composition to get the mean molecular weight, which enters the EOS, and the opacity profile, needed to calculate the luminosity term in Eq. (1) and the observed fluxes. In other words, we use the abundances at $t=t_{0}+\Delta t$ as a representation of the composition during the whole hydrodynamic step. This procedure produces all the information needed to evolve Eq. (1), and thus to update temperature and specific internal energy of the outflow.
@@
For matter that experienced hot NSE during the NR simulation, but underwent NSE freeze-out before reaching the extraction radius, the Boltzmann term in Eq. (8) tends to overestimate the abundances of the iron-group elements. Moreover, the impact of nuclear reactions happening immediately after NSE drop-out is neglected. To correct these effects, we implement a backtracking-NSE initialization similar to the one described in Radice et al. (2016); Perego et al. (2022). Essentially, if $T_{0}$, $\rho_{0}$, $Y_{e,0}$, $s_{0}$ are the properties at the extraction radius, we assume that the fluid element had expanded adiabatically from its NSE configuration at some $T_{\text{NSE}}>T_{0}$, $\rho_{\text{NSE}}$, $Y_{e,\text{NSE}}$, $s_{\text{NSE}}$. On the small timescales ($\lesssim 10$ ms) needed for the ejected matter to get to the extraction sphere, we expect the electron fraction and the entropy not to change significantly, so $Y_{e,\text{NSE}}\simeq Y_{e,0}$, $s_{\text{NSE}}\simeq s_{0}$. We check a posteriori that the relative changes in the entropy and electron fraction are on the order of 5% or less during the pre-evolution phase in most part of the ejecta (see Fig. 20). We finally estimate the density $\rho_{\text{NSE}}$ of the fluid element at NSE freeze-out by using the EOS employed in the NR simulation and starting from $T_{\text{NSE}}$, $Y_{e,\text{NSE}}=Y_{e,0}$, $s_{\text{NSE}}=s_{0}$. We calculate the correspondent NSE composition via Eq. (8), and then mimic the abundance evolution up to the extraction radius assuming homologous expansion. In particular, we parametrize the density evolution as in Eq. (1) of Lippuner & Roberts (2015), and estimate the expansion timescale $\tau_{\text{ex}}$ as in Radice et al. (2016) [^1] The prescription of Lippuner & Roberts (2015) is meant to approximate the physical evolution of a fluid element starting at early times after the ejection. The assumption of homologous expansion, not realistic at early times, is needed to get an estimate of the expansion timescale of the ejecta. Since variations of a factor of a few in $\tau_{\text{ex}}$ do not significantly affect the evolution of the NN (Lippuner & Roberts, 2015; Perego et al., 2021), our assumption gives a reliable approximation for the composition at the extraction radius.. To ensure that the evolved thermodynamic trajectory matches the physical conditions of the fluid element at the extraction radius, we also use an analytical prescription for the temperature history inspired by the adiabatic expansion law for an ideal gas, $T(\rho)=T_{\text{NSE}}\,(\rho(t)/\rho_{\text{NSE}})^{(\gamma-1)}$. The effective adiabatic factor $\gamma$, specific for each fluid element, is determined by $\gamma=1+\ln(\rho_{0}/\rho_{\text{NSE}})/\ln(T_{0}/T_{\text{NSE}})$, which ensures $T(\rho_{0})=T_{0}$. The composition at the end of the backtracking is used together with the originally extracted $s$ to define the initial condition of the fluid element.
@@
Low-entropy tidal ejecta (more efficiently produced for strongly asymmetric binaries) are expelled before the merger can raise the temperature, and are therefore expected to retain a composition close to the one set by the cold-NSE condition, for which $T\lesssim 1$ MeV. In principle, this would suggest using the cold-NSE method. However, in an NR simulation, numerical artifacts (e.g. shock heating propagating inward from the surface of the compact objects) can boost the initial temperature of the neutron stars up to few MeV. In all the NR simulations analyzed in this paper, the neutron stars reach this temperature in $\sim 1$ ms during the inspiral phase. The spurious high temperatures introduce small artifacts in the tracers calculations and in the initialized NSE composition, regardless of the method used. In Appendix B, we investigate the systematic uncertainties propagating from the different initialization methods to the nucleosynthesis and kilonova results.